OpenStudy (anonymous):

Help me simplify?? m^2n^4 ------- m^5n^3

OpenStudy (anonymous):

\[\frac{ m ^{2}n ^{4} }{ m ^{5}n ^{3} } = \frac{ n }{ m ^{3} }\]

OpenStudy (anonymous):

Thanks! Can you help me with this one too? x^4y^-1z^8 ----------- x^4y^-5z @Brigadier68

OpenStudy (anonymous):

\[\frac{ x ^{4}y _{-1}z ^{8} }{ x ^{4}y ^{-5}z }=y ^{4}z\]